# ToolRLA: Fine-Grained Reward Decomposition for Tool-Integrated Reinforcement Learning Alignment in Domain-Specific Agents

### v1

@@ Abstract
Tool-integrated reasoning agents that interleave natural language deliberation with external API calls have shown promise for complex, multi-step tasks. However, aligning such agents for deployment in high-stakes domain-specific settings remains challenging: existing reinforcement learning approaches rely on coarse binary rewards (success/failure) that provide insufficient guidance for the nuanced tool invocation behaviors required in production environments. We present ToolRLA, a three-stage post-training pipeline—Supervised Fine-Tuning (SFT), Group Relative Policy Optimization (GRPO), and Direct Preference Optimization (DPO)—designed for tool-integrated agents in domain-specific settings. The core contribution is a fine-grained reward function with multiplicative correctness decomposition that evaluates tool invocation along four orthogonal dimensions: format validity, tool selection correctness, invocation efficiency, and compliance with domain constraints. Within the correctness dimension, multiplicative composition encodes a natural priority ordering—correct tool selection is prerequisite to meaningful parameter evaluation—while a large negative compliance penalty () ensures regulatory adherence dominates task performance. We deploy ToolRLA on a real-world financial advisory copilot serving 80+ advisors with 1,200+ daily queries across 15+ heterogeneous backend APIs. After three months of production operation, ToolRLA achieves a 47% improvement in end-to-end task completion rate (62% 91%), a 63% reduction in tool invocation error rate (38% 14%), and a 93% reduction in regulatory violation rate (12% 0.8%), while maintaining sub-2-second end-to-end latency. Ablation studies confirm that fine-grained reward decomposition accounts for 7 percentage points of improvement over coarse additive reward design. We further validate generalizability on ToolBench and API-Bank public benchmarks.
ToolRLA: Fine-Grained Reward Decomposition for
Tool-Integrated Reinforcement Learning Alignment
in Domain-Specific Agents
@@ 1 Introduction
Second, reinforcement learning approaches for tool use typically employ coarse binary reward signals—a trajectory either succeeds or fails Patil et al. (2024); Du et al. (2024). Binary rewards provide insufficient gradient signal for the multi-dimensional quality requirements of domain-specific tool invocation: a trajectory that selects the correct tools but constructs malformed parameters is qualitatively different from one that selects the wrong tool entirely, yet both receive reward under binary evaluation. This coarseness slows convergence and fails to encode domain-specific priority orderings (e.g., regulatory compliance must dominate task completion).
@@
The central technical contribution is a fine-grained reward function that decomposes tool invocation quality along four orthogonal dimensions:
where is itself a multiplicative composition of tool name selection, required-tool coverage, and parameter accuracy scores, encoding a prerequisite ordering among correctness sub-dimensions. The multiplicative design encodes a natural precedence: correct tool selection is a prerequisite for meaningful parameter evaluation, and parameter correctness is prerequisite for a useful execution result. A large negative compliance term () ensures that regulatory violations are penalized more severely than complete task failure, establishing the priority ordering compliance correctness efficiency as an inductive bias in the reward landscape.
@@ 3.1 System Architecture: Single-Model ReAct Agent
Our production predecessor was a cascaded multi-model pipeline: intent classification slot filling hard-coded router execution model. While each module achieved 85–90% individual accuracy, the serial composition degraded end-to-end success on tasks requiring three or more steps to 62%. Moreover, the hard-coded router lacked any mechanism for mid-trajectory error recovery: once the router selected an incorrect branch, downstream modules received corrupted inputs with no opportunity to self-correct upon observing execution feedback.
We replaced this pipeline with a single-model ReAct agent that implements the Thought–Action–Observation loop:
@@
We expose 15 atomic tools and 5 composite tools, each specified as a four-tuple following the standard JSON Schema specification. Composite tools aggregate multiple atomic calls into a single invocation (e.g., GetClientOverview returns portfolio holdings, fund profiles, and recent transactions in one round-trip), reducing the average number of invocation rounds from 4.2 to 2.8. New tools can be integrated in 3–5 working days by (i) defining the JSON Schema, (ii) writing a thin adapter for format normalization, and (iii) supplementing 50–100 targeted training examples followed by a brief GRPO fine-tuning pass.
@@ 3.3.2 Fine-Grained Reward Function
The total reward decomposes additively across four dimensions:
@@ 5.3 Public Benchmark Results
ToolRLA achieves 51.3% Pass Rate on ToolBench, surpassing the GPT-4 function-calling baseline (46.2%) by 5.1pp. On API-Bank it reaches 71.8% Call Accuracy, a 4.7pp gain over the same baseline. AvaTaR’s contrastive trajectory reasoning (44.3%) validates that contrasting successful and failed trajectories yields meaningful signal—yet falls 7pp short of ToolRLA, reflecting the advantage of RL-shaped parameter-level feedback over prompt optimization alone. The joint efficiency scheduling of Bloomberg AI Engineering (2025) (48.2%) narrows the gap further, confirming that structured tool-use training is competitive with frontier closed-source models. The remaining 3.1pp advantage of ToolRLA over the ACL 2025 baseline is consistent with the 7pp multiplicative-vs-additive gap in our internal ablation, suggesting that fine-grained reward decomposition provides a generalizable inductive bias across API ecosystems.
@@ 5.4 Online Deployment Results
The 67% reduction in manual retry rate (28% 9%) and the 60% reduction in abandonment rate (35% 14%) reflect the agent’s improved reliability in daily advisory workflows. Daily query volume grew 50% (800 1,200+), partly driven by advisor confidence in the system’s responses. Satisfaction scores were collected via bi-weekly surveys across four dimensions (answer accuracy, response speed, tool reliability, compliance safety), controlling for novelty effects by tracking the score trajectory: week 1 (4.0) week 4 (4.2) week 12 (4.3, stable), confirming that improvements are not attributable to initial adoption enthusiasm.
@@ 6.3 Reward Signal Dynamics
The fraction of group-8 samples with rises from 45% (SFT initialization) to 78% by the end of GRPO training, confirming the reward signal remains non-degenerate throughout. triggers on 3% of trajectories after 1,000 steps; the residual grey-area violations motivate the subsequent DPO stage.
@@ 6.4 Model Size and Training Efficiency
Qwen3-14B closes 84% of the TCR gap between 7B and 32B while maintaining sub-2s latency—the threshold acceptable for real-time advisory workflows. The 32B variant provides only marginal gains (+2.5pp TCR, 1.8pp TIER) at 2.4 higher inference cost, making it impractical under our latency budget. GRPO convergence stabilizes at 8,000 training steps: the fraction of group-8 samples with positive advantage exceeds 70% by step 4,000 and plateaus at 78% by step 8,000, with no measurable gain beyond 10,000 steps (0.5pp TIER on held-out validation). This fast convergence—relative to typical PPO schedules requiring 50K rollouts for comparable gains—is attributable to the fine-grained reward signal reducing reward sparsity and providing dense per-dimension gradient information even for partially correct trajectories.
@@ 8 Conclusion
We presented ToolRLA, a three-stage post-training framework for tool-integrated agents in domain-specific settings. The central contribution is a fine-grained, multiplicatively decomposed reward function that decomposes tool invocation quality along four dimensions and encodes task-specific priority orderings as inductive biases in the reward landscape. Ablation studies demonstrate that multiplicative composition of the correctness reward accounts for 7 percentage points of TIER improvement over additive alternatives, and that the three-stage pipeline (SFT GRPO DPO) is strictly better than any prefix thereof. Deployed on a production financial advisory copilot over three months, ToolRLA delivers a 47% improvement in task completion rate, a 63% reduction in tool invocation errors, and a 93% reduction in regulatory violations. We hope these results encourage the community to move beyond binary reward signals toward structured, semantics-aware reward decomposition for tool-integrated reinforcement learning.

### v3

# ToolRLA: Multiplicative Reward Decomposition for Tool-Integrated Agents

@@ Abstract
Tool-integrated agents that interleave reasoning with API calls are promising for complex tasks, yet aligning them for high-stakes, domain-specific deployment remains challenging: existing reinforcement learning approaches rely on coarse binary rewards that cannot distinguish tool selection errors from malformed parameters. We present ToolRLA, a three-stage post-training pipeline (SFT GRPO DPO) for domain-specific tool agents. The core contribution is a fine-grained reward function with multiplicative correctness decomposition spanning four dimensions—format validity, tool selection, parameter accuracy, and regulatory compliance—that encodes domain priority orderings as inductive biases in the reward landscape. Deployed on a financial advisory copilot (80+ advisors, 1,200+ daily queries), ToolRLA achieves over three months: a 47% improvement in task completion rate (62%91%), a 63% reduction in tool invocation errors (38%14%), and a 93% reduction in regulatory violations (12%0.8%), within sub-2-second latency. Ablation studies show the multiplicative reward design accounts for 7 percentage points of improvement over additive alternatives. Generalization is further validated on ToolBench and API-Bank.
ToolRLA: Multiplicative Reward Decomposition for Tool-Integrated Agents
@@ 1 Introduction
Second, reinforcement learning approaches for tool use typically employ coarse binary reward signals—a trajectory either succeeds or fails Patil et al. (2024); Du et al. (2024). Binary rewards provide insufficient gradient signal for the multi-dimensional quality requirements of domain-specific tool invocation: a trajectory that selects the correct tools but constructs malformed parameters is qualitatively different from one that selects the wrong tool entirely, yet both receive reward under binary evaluation. This coarseness slows convergence and fails to encode domain-specific priority orderings (e.g., regulatory compliance must dominate task completion). Figure 1 illustrates this limitation and motivates our fine-grained decomposition.
@@
The central contribution is a fine-grained reward function decomposed along four dimensions: format (), correctness (), efficiency (), and compliance (). Critically, is a multiplicative composition of tool-name, coverage, and parameter accuracy sub-scores, so a wrong tool selection collapses correctness regardless of parameter quality. A large negative compliance penalty () enforces compliance correctness efficiency as an inductive bias in the reward landscape.
@@ 3.1 System Architecture: Single-Model ReAct Agent
Our production predecessor was a cascaded multi-model pipeline (intent classifier slot filler router), which degraded end-to-end success to 62% and lacked mid-trajectory error recovery. We replaced it with a single-model ReAct agent that implements the Thought–Action–Observation loop:
@@
We expose 15 atomic tools and 5 composite tools, each specified as a four-tuple following the standard JSON Schema specification. Composite tools aggregate multiple atomic calls into a single invocation (e.g., GetClientOverview returns portfolio holdings, fund profiles, and recent transactions in one round-trip), reducing average invocation rounds from 4.2 to 2.8.
@@ 3.3.2 Fine-Grained Reward Function
The total reward decomposes additively across four dimensions (Figure 3):
@@ 5.3 Public Benchmark Results
ToolRLA achieves 51.3% Pass Rate on ToolBench (+5.1pp over GPT-4 function calling) and 71.8% Call Accuracy on API-Bank (+4.7pp). AvaTaR’s contrastive trajectory reasoning (44.3%) falls 7pp short, reflecting the advantage of RL-shaped parameter-level feedback over prompt optimization alone. The 3.1pp advantage over Bloomberg AI Engineering (2025) (48.2%) is consistent with the 7pp multiplicative-vs-additive gap in our ablation, suggesting fine-grained reward decomposition provides a generalizable inductive bias across API ecosystems.
@@ 5.4 Online Deployment Results
The 67% reduction in manual retry rate (28%9%) and 60% reduction in abandonment rate (35%14%) reflect improved reliability in daily advisory workflows. Daily query volume grew 50% (8001,200+), driven by advisor confidence in the system’s responses. Satisfaction scores stabilized at 4.3/5 by week 12 (up from 3.1), ruling out novelty effects.
@@ 6.3 Reward Signal Dynamics
Figure 4 plots the reward signal dynamics throughout GRPO training. The fraction of group-8 samples with rises from 45% (SFT initialization) to 78% by convergence, confirming the reward signal remains non-degenerate throughout. triggers on 3% of trajectories after 1,000 steps; the residual grey-area violations motivate the subsequent DPO stage.
@@ 6.4 Model Size and Training Efficiency
Qwen3-14B closes 84% of the TCR gap between 7B and 32B while maintaining sub-2s latency. The 32B variant provides only marginal gains (+2.5pp TCR, 1.8pp TIER) at 2.4 higher inference cost, making it impractical under our latency budget. GRPO convergence stabilizes at 8,000 steps (78% of samples with positive advantage), faster than typical PPO schedules requiring 50K rollouts, attributable to the fine-grained reward reducing sparsity by providing dense per-dimension gradient signal even for partially correct trajectories.
@@ 8 Conclusion
We presented ToolRLA, a three-stage post-training framework for tool-integrated agents in domain-specific settings. The central contribution is a fine-grained multiplicative reward function that evaluates tool invocation quality along four dimensions and encodes task-specific priority orderings as inductive biases in the reward landscape. Ablation studies demonstrate that multiplicative composition of the correctness reward accounts for 7 percentage points of TIER improvement over additive alternatives, and that the three-stage pipeline (SFT GRPO DPO) is strictly better than any prefix thereof. Deployed on a production financial advisory copilot over three months, ToolRLA delivers a 47% improvement in task completion rate, a 63% reduction in tool invocation errors, and a 93% reduction in regulatory violations. These results establish structured, semantics-aware reward decomposition as a practically effective direction for tool-integrated reinforcement learning beyond binary feedback signals.